solving linear systems using substitution, what is
4x-7y=10
y=x-7

Answer by mananth(16946)

4x-7y=10 .............1
x-y=7.............2

multiply (1)by -1
Multiply (2) by 7
-4x+7y=-10
7x-7y=49

Add the two equations

3x=39

/3

x=13

plug value of x in (1)

4x-7y=10
52-7y=10
-7y=10-52
-7y=-42
y=6